How many gallons of pure alcohol should be mixed with 20 gal of 15% alcohol solution to obtain a mixture that is 25% alcohol?
Answer by jorel1380(3719)

x+20(.15)=(x+20).25
x+3=.25x+5
.75x=2
x=2/.75=2.67
You must mix 2.67 gals. of pure alcohol to obtain the desired mixture..